OpenStudy (anonymous):

1 2 3 4 5 6 7 8 9 10 . Solve the rational equation. 5 x - 1 − 3 x + 2 = 0

OpenStudy (ash2326):

Is this your question \[\frac{5}{x-1}-\frac{3}{x+2}=0\]